finding this variable a,b,c and d
a+b+c+d=0
-a+b+2c+3d=0
-a+b+8c+27d=0
1
Expert's answer
2019-05-14T13:15:30-0400

a + b + c + d = 0,


- a + b + 2c + 3d = 0,


- a + b + 8c + 27d = 0,


a = b + 8c + 27d,


a = b + 2c + 3d,


b + 8c + 27d = b + 2c + 3d,


8c + 27d = 2c + 3d,


24d = - 6c,


-4d = c


a = - b - c - d,


b = - 2c - 3d + a = 8d - 3d + (- b - (-4d) - d) = 8d - b,


2b = 8d,


b = 4d,


a = - 4d + 4d -d = -d